log base1/3 p < 0
1. Is there a sign flip?
2. Well, the base is less than one and now what do I do?

Answer by lwsshak3(11628)

log base1/3 p < 0
1. Is there a sign flip?
2. Well, the base is less than one and now what do I do?
***
in exponential form, this reads: base(1/3) raised to negative number=p
say, use -1 as the negative number:
(1/3)^-1=p
1/(1/3)=p
3=p
yes, there is a sign flip
I think